# Quadrature Coherence Scale of Linear Combinations of Gaussian Functions in Phase Space

### v1

@@ Abstract
The quadrature coherence scale is a recently introduced measure that was shown to be an efficient witness of nonclassicality. It takes a simple form for pure and Gaussian states, but a general expression for mixed states tends to be prohibitively unwieldy. In this paper, we introduce a method for computing the quadrature coherence scale of quantum states characterized by Wigner functions expressible as linear combinations of Gaussian functions. Notable examples within this framework include cat states, GKP states, and states resulting from Gaussian transformations, measurements, and breeding protocols. In particular, we show that the quadrature coherence scale serves as a valuable tool for examining the scalability of nonclassicality in the presence of loss. Our findings lead us to put forth a conjecture suggesting that, subject to 50% loss or more, all pure states become classical. We also consider the quadrature coherence scale as a measure of quality of the output state of the breeding protocol.
1 Introduction
@@
Counterintuitive quantum phenomena including superposition and entanglement have transformed from problems [1] to curiosities [2, 3, 4, 5] to features that confer advantages in multiple domains [6, 7, 8, 9, 10, 11, 12, 13]. Superposition and entanglement are based off of coherence, which underlies all interference effects. This makes the generation [14, 15, 16], manipulation [17, 18, 19, 20], quantification [21, 22, 23, 24, 25, 26, 27], and measurement [28, 29, 30, 31, 32, 33] of coherence essential tasks, especially when viewed through the lens of coherence as a resource [21, 22, 34, 35, 36].
Coherence between macroscopically distinct states is a good indicator of quantumness or nonclassicality. The quadrature coherence scale (QCS) was recently introduced as such an indicator, quantifying both the amount of coherence and the macroscopicity of the coherence in a quantum state; the QCS is a bona fide witness of optical nonclassicality [37, 38]. An interferometric scheme with two identical copies of a state suffices for measuring the QCS of the state without requiring full state tomography [39], as was recently demonstrated on a cloud quantum computer [40].
@@
For states that are pure or Gaussian, the QCS is simple to compute. However, a general computation requires calculations involving multiple integrals over a state’s entire Wigner function, which is essential to practical situations because non-Gaussian states are crucial to linear optical quantum computation [41, 42, 43, 44] and pure states are rare.
@@
In this work, we show that computation of the QCS is significantly easier for states that can be written as superpositions and convex combinations of Gaussian states, which includes non-Gaussian states that are important to quantum metrology and computation. This takes advantage of a recently developed formalism for states whose Wigner functions are linear combinations of not-necessarily-real Gaussian functions [45] and for simulating quantum optics with finite-rank superpositions of coherent states [46].
@@
As an example, Schrödinger cat states are important non-Gaussian states used for demonstrating quantumness [47], quantum-enhanced measurement sensitivity [48], and error-correcting codes [49] and belong to this class of superpositions of Gaussian states. They can be combined via linear optical networks and homodyne measurement can be performed on some branches while remaining in this class, allowing them to iteratively “breed” [50, 51, 15] into another important type of states known as Gottesman-Kitaev-Preskill (GKP) states [52]. Moreover, such states subject to the dominant noise source for photonics, viz., loss, still remain in this class. These allow the QCS to measure the nonclassicality and thereby the quality of a state intended for fault-tolerant quantum computation as it improves via breeding and degrades via loss.
@@
Our framework is fully general and allows the computation of the QCS of any state belonging to the correct class subject to any Gaussian transformations and measurements. It was developed with GKP states in mind and may prove useful throughout multimode quantum optics with continuous variables.
@@
The paper is divided as follows. In the proceeding two sections, we start by introducing the main concept of quantum optics in phase space and then introduce the quadrature coherence scale. In Sec. 4, we define the formalism related to the expression of the Wigner function of a state as a linear combination of Gaussian functions and in Sec. 5 we show how to compute the QCS for such states. In Sec. 6, we give concrete examples: we compute the QCS of cat states and GKP states, we study the scalability of the QCS through a loss channel, and we consider the QCS as a measure of fidelity for the output of a breeding protocol. We finally conclude in Sec. 7 and discuss some future research problems.
2 Phase space formalism
@@
In this section, we provide a brief overview of the symplectic formalism employed for continuous-variable states in quantum optics. More details can be found, for example, in [53, 54].
@@
3 Computation of the QCS – General
A state $\rho$ is said to be (optically) classical [55] if and only if there exists a positive function $P(\alpha)$, called the Sudarshan-Glauber P function, such that
@@ 3 Computation of the QCS – General
where $|\alpha\rangle=D(\alpha)|0\rangle$ is a coherent state, $D(\alpha)$ is the displacement operator, and $|0\rangle$ is the vacuum state. Otherwise, the state is said to be nonclassical. In other words, a state is said to be nonclassical if it is not a probabilistic mixture of coherent states. It is typically challenging to determine if such a P-function exists and even more so to assess its positivity. Therefore, there is a need for defining measures and witnesses of nonclassicality. The most famous one is probably the negativity of the Wigner function: if the Wigner function of a state $\rho$ takes negative values at any point on the phase space, the state is nonclassical [56]. Nevertheless, many other measures and witness exist [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80]. In this work, we focus on the quadrature coherence scale (QCS), a recently introduced witness of nonclassicality [38, 37, 81, 40, 82].
Nonclassicality is linked to the coherences present in a state, but, while many nonclassicality witnesses and measures consider the size of the coherences, the QCS also measures where those coherences are located, showcasing a different aspect of the nonclassicality of the state. As an example, the nonclassicality of a squeezed state will be detected by the QCS while it is concealed from the Wigner negativity. The QCS was calculated for several benchmark states including Fock states, squeezed thermal states, and cat states [37], as well as states with more complicated QCSs like photon-added and subtracted Gaussian states [82].
@@
For a pure state $|\psi\rangle$, it simplifies to the total noise (i.e. the sum of the variances). It is thus the trace of the covariance matrix $\boldsymbol{\gamma}$:
@@
The QCS is not a measure, but a witness of nonclassicality. All classical states have ${\mathcal{C}}_{\rho}\leq 1$. Therefore, a value of the QCS greater than 1 certifies that the state is nonclassical. In addition, despite not being a proper measure of nonclassicality, it was proven in [38] that the distance D($\rho$,$\mathcal{E}_{cl}$) between the state $\rho$ and the set of nonclassical states $\mathcal{E}_{cl}$ is bounded by the QCS in the following way:
@@
4 Linear combination of Gaussian functions in phase space
@@
Since a Wigner function is normalized to $1$, we have $\sum_{m}c_{m}=1$. Note that, if $G_{\boldsymbol{\mu}_{m},\boldsymbol{\gamma}_{m}}$ represents a valid quantum state, then it is equivalent to Eq. (3) and is real. However, $G_{\boldsymbol{\mu}_{m},\boldsymbol{\gamma}_{m}}(x)$ can also be a complex (normalized) function, with Eq. (11) still describing a valid quantum state. More details can be found in Ref. [45] where this definition was introduced.
Writing the Wigner function in the form of Eq. (11) is useful to describe, for example, cat states, Fock states, and GKP states. We will develop this last example in the next section. Even more interesting is that this formalism is useful in describing the resultant state when a Gaussian transformation or measurement is applied to a state of the form of Eq. (11). Indeed, a Gaussian transformation maps a Gaussian state into another Gaussian state; hence by linearity, a state of the form Eq. (11) will still be written in this specific form after the Gaussian transformation. In addition, a deterministic Gaussian map (such a loss channel) will change neither the number nor the weights of the coefficients $c_{m}$ in the sum, and each covariance matrix $\boldsymbol{\gamma}_{m}$ and mean value vector $\boldsymbol{\mu}_{m}$ will be updated according to Eq. (4). Once again, we refer to Ref. [45] for more details regarding the application of Gaussian transformations to states written as a linear combination of Gaussian functions. We also draw the reader’s attention to Ref. [83] for a more advanced work with linear combinations of Gaussian functions.
Note that, in theory, every state can be formatted in this way, but the description might involve arbitrary infinite sums or integrals that makes this formalism impractical.
5 Computation of the QCS for a state whose Wigner function can be written as a sum of Gaussian functions
The denominator of Eq. (7) is proportional to the purity $\mathrm{Tr}\rho^{2}$ of the state, which can be computed as follows:
@@
where we used that the product of two Gaussian functions gives a new Gaussian function (see for example Eq. (337) in Ref. [84]), with
@@
Since the Gaussian function is normalized, $\int G_{\boldsymbol{d},\boldsymbol{\Gamma}}(\boldsymbol{r})\mathrm{d}^{2}\boldsymbol{r}=1$. In addition, we can write $\boldsymbol{\Gamma}=\boldsymbol{\gamma}_{m}(\boldsymbol{\gamma}_{m}+\bar{\boldsymbol{\gamma}}_{n})^{-1}\bar{\boldsymbol{\gamma}}_{n}$ [84], which implies $\det\boldsymbol{\Gamma}=\det(\boldsymbol{\gamma}_{m})\det(\bar{\boldsymbol{\gamma}}_{n})/\det(\boldsymbol{\gamma}_{m}+\bar{\boldsymbol{\gamma}}_{n})$. Thus,
@@
In a similar way, the numerator of Eq. (7) can be computed as follows:
@@
Here, $E[x]$ is the mean value of $x$ with respect to the probability distribution $G_{d,\boldsymbol{\gamma}}$. With Eq. (357) in Ref. [84] and a bit of algebra, we have
@@
6 Examples
6.1 Cat states
Let us start with the simple example of a cat state defined as $|cat\rangle=\sqrt{\mathcal{N}}(|\alpha\rangle+|-\alpha\rangle)$ where $|\alpha\rangle$ is a coherent state and $\mathcal{N}=(2+2{\mathrm{e}}^{-2|\alpha|^{2}})^{-1}$ a normalization constant. The larger the value of $\alpha$, the more macroscopically distinguishable the two terms in the coherent superposition become and thus the larger we expect the QCS to be. Since it is a pure state, the QCS of a cat state can be computed as the sum of its variances (see Eq. (8)) and we find [37]
@@ 6.1 Cat states
To use our method, we write the cat state as the sum of four Gaussian functions with the following parameters and coefficients [45]:
@@
6.2 GKP states
Let us now consider a more interesting example, that is the grid state introduced by Gottesman-Kitaev-Preskill (GKP) [52]. An ideal GKP state is represented by an infinite number of Dirac delta functions equally spaced by $2\sqrt{\pi}$ in the phase space:

### v3

@@ Abstract
The quadrature coherence scale (QCS) is a recently introduced measure that was shown to be an efficient witness of nonclassicality. It takes a simple form for pure and Gaussian states, but a general expression for mixed states tends to be prohibitively unwieldy. In this paper, we introduce a method for computing the quadrature coherence scale of quantum states characterized by Wigner functions expressible as linear combinations of Gaussian functions. Notable examples within this framework include cat states, GKP states, and states resulting from Gaussian transformations, measurements, and breeding protocols. In particular, we show that the quadrature coherence scale serves as a valuable tool for examining the scalability of nonclassicality in the presence of loss. Our findings lead us to put forth a conjecture suggesting that, subject to 50% loss or more, all pure states lose any QCS-certifiable nonclassicality. We also consider the quadrature coherence scale as a measure of quality of the output state of the breeding protocol.
I Introduction
@@
Counterintuitive quantum phenomena including superposition and entanglement have transformed from problems EPR1935 to curiosities Bell1964 ; Bell1966 ; KochenSpecker1967 ; vanFraassen1982 to features that confer advantages in multiple domains Ekert1991 ; Bennettetal1993 ; Shor1995 ; Grover1996 ; Giovannettietal2004 ; Dowling2008 ; Aberg2014 ; Korzekwaetal2016 . Superposition and entanglement are based off of coherence, which underlies all interference effects. This makes the generation Breeding_Etesse_2 ; Breeding ; GoldbergSteinberg2020 , manipulation ManiKarimipour2015 ; Bromleyetal2015 ; GarciaDiazetal2016 ; Buetal2017 , quantification Aberg2006arxiv ; Baumgratzetal2014 ; Streltsovetal2015 ; Ranaetal2016 ; Yuetal2016 ; Rastegin2016 ; Hillery2016 , and measurement Girolami2014 ; Wangetal2017 ; Zhengetal2018 ; Xuetal2020 ; Wuetal2021 ; Yuanetal2023 of coherence essential tasks, especially when viewed through the lens of coherence as a resource Aberg2006arxiv ; Baumgratzetal2014 ; WinterYang2016 ; Streltsovetal2017 ; Liuetal2020 .
Coherence between macroscopically distinct states is a good indicator of quantumness or nonclassicality. Several measures and witnesses of nonclassicality have been introduced, with their strengths and weaknesses. For example, the distance to the set of classical set Hillery1 is a precise measure of non-classicality, but is typically difficult to evaluate because it requires an optimization over a set of states. The negative volume of the Wigner function Kenfack is also a very popular witness that can be hard to compute due to looking for small, rapid interferences via numerical integration. The latter also fails to detect many nonclassical states; e.g., squeezed states. The quadrature coherence scale (QCS) was recently introduced as such an indicator, quantifying both the amount of coherence and the macroscopicity of the coherence in a quantum state; the QCS is a bona fide witness of optical nonclassicality Hertz ; Debievre with the important advantage of bounding the distance to the set of classical states while being relatively easy to compute. Let us note also that for pure states QCS is equivalent to many other nonclassicality witnesses Goldbergetal23 . In addition, an interferometric scheme with two identical copies of a state suffices for measuring the QCS of the state without requiring full state tomography Griffet , unlike the nonclassicality distance or Wigner negativity that require full tomography and then full optimization or integration; the QCS measurement was recently demonstrated on a cloud quantum computer Goldbergetal23 .
@@
For states that are pure or Gaussian, the QCS is simple to compute. However, a general computation requires calculations involving multiple integrals over a state’s entire Wigner function, which is essential to practical situations because non-Gaussian states are crucial to linear optical quantum computation LloydBraunstein1999 ; Bartlettetal2002 ; MariEisert2012 ; Walschaers21 and pure states are rare.
@@
In this work, we show that computation of the QCS is significantly easier for states that can be written as superpositions and convex combinations of Gaussian states, which includes non-Gaussian states that are important to quantum metrology and computation. This takes advantage of a recently developed formalism for states whose Wigner functions are linear combinations of not-necessarily-real Gaussian functions XanaduSumGaussian and for simulating quantum optics with finite-rank superpositions of coherent states MarshallAnand2023 .
@@
This description of states might sound restrictive, but it actually represents a large family of relevant states. As an example, Schrödinger cat states are important non-Gaussian states used for demonstrating quantumness Schrodinger1935cat , quantum-enhanced measurement sensitivity Zurek2001 , and error-correcting codes Cochraneetal1999 and belong to this class of superpositions of Gaussian states. They can be combined via linear optical networks and homodyne measurement can be performed on some branches while remaining in this class, allowing them to iteratively “breed” Breeding_Vasconcelos ; TerhalWeigand2016 ; Breeding into another important type of states known as Gottesman-Kitaev-Preskill (GKP) states GKPpaper . Moreover, such states subject to the dominant noise source for photonics, viz., loss, still remain in this class, as do states subject to any Gaussian channels. In contrast, other nonclassicality measures such as quantum Fisher information that may be easy to compute for pure states must be computed from scratch with more challenging calculations when the state becomes mixed due to loss. These allow the QCS to measure the nonclassicality and thereby the quality of a state intended for fault-tolerant quantum computation as it improves via breeding and degrades via loss.
@@
In addition, using the description of the P-function (see eq. (6)), one can see that any state can be written as a combination of Gaussian (coherent) states, at least formally, because the coefficients in this combination need not be positive. Moreover, it was recently proven that any pure state can be represented with arbitrary precision as a finite combination of coherent states, while all mixed states are just convex combinations of such MarshallAnand2023 . Our framework is thus fully general and allows the computation of the QCS of any state belonging to the correct class subject to any Gaussian transformations and measurements. It was developed with GKP states in mind and may prove useful throughout multimode quantum optics with continuous variables.
@@
The paper is divided as follows. In the proceeding two sections, we start by introducing the main concept of quantum optics in phase space and then introduce the quadrature coherence scale. In Sec. IV, we define the formalism related to the expression of the Wigner function of a state as a linear combination of Gaussian functions and in Sec. V we show how to compute the QCS for such states. In Sec. VI, we give concrete examples: we compute the QCS of cat states and GKP states, we study the scalability of the QCS through a loss channel, and we consider the QCS as a measure of fidelity for the output of a breeding protocol. In Sec. VII we discuss other nonclassicality measures. We finally conclude in Sec. VIII and discuss some future research problems.
II Phase space formalism
@@
In this section, we provide a brief overview of the symplectic formalism employed for continuous-variable states in quantum optics. More details can be found, for example, in weedbrook ; Anaellethesis .
@@
III Computation of the QCS – General
A state $\rho$ is said to be (optically) classical Titulaer if and only if there exists a positive function $P(\alpha)$, called the Sudarshan-Glauber P function, such that
@@ III Computation of the QCS – General
where $|\alpha\rangle=D(\alpha)|0\rangle$ is a coherent state, $D(\alpha)$ is the displacement operator, and $|0\rangle$ is the vacuum state. Otherwise, the state is said to be nonclassical. In other words, a state is said to be nonclassical if it is not a probabilistic mixture of coherent states.
It is typically challenging to determine if such a P-function exists and even more so to assess its positivity. Therefore, there is a need for defining measures and witnesses of nonclassicality. The most famous one is probably the negativity of the Wigner function: if the Wigner function of a state $\rho$ takes negative values at any point on the phase space, the state is nonclassical Kenfack . Nevertheless, many other measures and witness exist Hillery1 ; Bach ; Hillery3 ; Lee ; Agarwal2 ; Lutkenhaus ; Dodonov ; Marian ; Richter ; Asboth ; Ryl ; Sperling ; Killoran ; Alexanian ; Nair ; Ryl2 ; Yadin ; Kwon2 ; Takagi18 ; Horoshko ; Luo ; Bohmann ; Tan2020 . In this work, we focus on the quadrature coherence scale (QCS), a recently introduced witness of nonclassicality Debievre ; Hertz ; HertzCerfDebievre ; Goldbergetal23 ; QCS_photon_added .
Nonclassicality is linked to the coherences present in a state, but, while many nonclassicality witnesses and measures consider the size of the coherences, the QCS also measures where those coherences are located, showcasing a different aspect of the nonclassicality of the state. As an example, the nonclassicality of a squeezed state will be detected by the QCS while it is concealed from the Wigner negativity.
The QCS was calculated for several benchmark states including Fock states, squeezed thermal states, and cat states Hertz , as well as states with more complicated QCSs like photon-added and subtracted Gaussian states QCS_photon_added .
@@
For a pure state $|\psi\rangle$, it simplifies to the total noise (i.e. the sum of the variances) and many other indicators of nonclassicality as enumerated in Ref. Goldbergetal23 . It is thus the trace of the covariance matrix $\boldsymbol{\gamma}$:
@@
The QCS is not a measure, but a witness of nonclassicality. All classical states have ${\mathcal{C}}_{\rho}\leq 1$. Therefore, a value of the QCS greater than 1 certifies that the state is nonclassical. In addition, despite not being a proper measure of nonclassicality, it was proven in Debievre that the distance [^3] The distance is defined as D($\rho$,$\mathcal{E}_{cl})=\rm{inf}_{\sigma\in\mathcal{E}_{cl}}|||\tilde{\rho}-\tilde{\sigma}|||$ with $|||A|||=\sqrt{\langle A|A\rangle}$ and $\tilde{\rho}=\rho/\sqrt{\mathrm{Tr}\rho^{2}}$. D($\rho$,$\mathcal{E}_{cl}$) between the state $\rho$ and the set of nonclassical states $\mathcal{E}_{cl}$ is bounded by the QCS in the following way:
@@
IV Linear combination of Gaussian functions in phase space
@@
Since a Wigner function is normalized to $1$, we have $\sum_{m}c_{m}=1$. Note that, if $G_{\boldsymbol{\mu}_{m},\boldsymbol{\gamma}_{m}}$ represents a valid quantum state, then it is equivalent to Eq. (3) and is real. However, $G_{\boldsymbol{\mu}_{m},\boldsymbol{\gamma}_{m}}(x)$ can also be a complex (normalized) function, with Eq. (11) still describing a valid quantum state. More details can be found in Ref. XanaduSumGaussian where this definition was introduced.
Writing the Wigner function in the form of Eq. (11) is useful to describe, for example, cat states, Fock states, and GKP states. We will develop this last example in the next section. Even more interesting is that this formalism is useful in describing the resultant state when a Gaussian transformation or measurement is applied to a state of the form of Eq. (11). Indeed, a Gaussian transformation maps a Gaussian state into another Gaussian state; hence by linearity, a state of the form Eq. (11) will still be written in this specific form after the Gaussian transformation. In addition, a deterministic Gaussian map (such a loss channel) will change neither the number nor the weights of the coefficients $c_{m}$ in the sum, and each covariance matrix $\boldsymbol{\gamma}_{m}$ and mean value vector $\boldsymbol{\mu}_{m}$ will be updated according to Eq. (4). Once again, we refer to Ref. XanaduSumGaussian for more details regarding the application of Gaussian transformations to states written as a linear combination of Gaussian functions. We also draw the reader’s attention to Ref. yao2023design for a more advanced work with linear combinations of Gaussian functions.
Note that, in theory, every state can be formatted in this way, but the description might involve arbitrary infinite sums or integrals that makes this formalism impractical. If one is willing to tolerate any arbitrarily small error in a description of their state, then any state can be written in this form with only a finite number of terms MarshallAnand2023 .
V Computation of the QCS for a state whose Wigner function can be written as a sum of Gaussian functions
The crucial observation is that all of the integrals required for computing the QCS simplify to sums of integrals of products of Gaussians. Since products of Gaussians yield new Gaussians and integration of Gaussians is well understood, closed-form solutions for the QCS of this class of states can be established, as follows.
The denominator of Eq. (7) is proportional to the purity $\mathrm{Tr}\rho^{2}$ of the state, which can be computed as follows, using def. (11) of the Wigner function:
@@
where we used that the product of two Gaussian functions gives a new Gaussian function (see for example Eq. (337) in Ref. MatrixCookBook ), with
@@
Since the Gaussian function is normalized, $\int G_{\boldsymbol{d},\boldsymbol{\Gamma}}(\boldsymbol{r})\mathrm{d}^{2}\boldsymbol{r}=1$. In addition, we can write $\boldsymbol{\Gamma}=\boldsymbol{\gamma}_{m}(\boldsymbol{\gamma}_{m}+\bar{\boldsymbol{\gamma}}_{n})^{-1}\bar{\boldsymbol{\gamma}}_{n}$ MatrixCookBook , which implies $\det\boldsymbol{\Gamma}=\det(\boldsymbol{\gamma}_{m})\det(\bar{\boldsymbol{\gamma}}_{n})/\det(\boldsymbol{\gamma}_{m}+\bar{\boldsymbol{\gamma}}_{n})$. Thus,
@@
In a similar way, noting that the gradient of a Gaussian is given by $\nabla G_{\boldsymbol{\mu}_{m},\boldsymbol{\gamma}_{m}}(\boldsymbol{r})=\boldsymbol{\gamma}_{m}^{-1}\left(\boldsymbol{r}-\boldsymbol{\mu}_{m}\right)G_{\boldsymbol{\mu}_{m},\boldsymbol{\gamma}_{m}}(\boldsymbol{r})$,the numerator of Eq. (7) can be computed as follows:
@@
Here, $E[x]$ is the mean value of $x$ with respect to the probability distribution $G_{d,\boldsymbol{\gamma}}$. With Eq. (357) in Ref. MatrixCookBook and a bit of algebra, we have
@@
VI Examples
VI.1 Cat states
Let us start with the simple example of a cat state defined as $|cat\rangle=\sqrt{\mathcal{N}}(|\alpha\rangle+|-\alpha\rangle)$ where $|\alpha\rangle$ is a coherent state and $\mathcal{N}=(2+2{\mathrm{e}}^{-2|\alpha|^{2}})^{-1}$ a normalization constant. The larger the value of $\alpha$, the more macroscopically distinguishable the two terms in the coherent superposition become and thus the larger we expect the QCS to be. Since it is a pure state, the QCS of a cat state can be computed as the sum of its variances (see Eq. (8)) and we find Hertz
@@ VI.1 Cat states
To use our method, we write the cat state as the sum of four Gaussian functions with the following parameters and coefficients XanaduSumGaussian :
@@
VI.2 GKP states
Let us now consider a more interesting example, that is the grid state introduced by Gottesman-Kitaev-Preskill (GKP) GKPpaper . An ideal GKP state is represented by an infinite number of Dirac delta functions equally spaced by $2\sqrt{\pi}$ in the phase space: